OpenStudy (anonymous):

[6.02] For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8

OpenStudy (zehanz):

You can do the following with equation 2: Add x on both sides, to get -y = -8 + x If you then add 8 on both sides, you have isolated x and you'll know what to substitute in the first eqn.

OpenStudy (anonymous):

-8+Y?

OpenStudy (zehanz):

No, because if you have -y = -8 + x, which is of course the same as: -8 +x = -y, if you ADD 8 on both sides, on the left side, you have -8 + 8 + x = x And on the right side, you have -y + 8. So in the end it is x = -y + 8.